An object is projected horizontally at 15 m/s. At the same time, another object located 9 meters down range is dropped from rest. When will the two objects collide?
1
Expert's answer
2018-09-19T15:40:08-0400
y_1=y_2=-(gt^2)/2.
x=vt
t=x/v=9/15=0.6 s.